Need help ASAP
Problem 2-13 Market Values and Book Values [LO 1)
Answer #1

Solution :-

(A) Book value of net fixed assets = $4.6 million = $4600000

Net working capital =

Current assets - Current Liabilities = 148000

Current Assets - 890000 = 148000

Current assets = 890000 + 148000 = $1038000

Therefore book value of total assets = net fixed assets + current assets = 4600000 + 1038000 = $5638000

(B)

Maket value of net working capital = Market value of Current assets - Current liabilities = 1005000 - 890000 = 115000$

Market value of Fixed assets = $8 million = $8000000

Therefore total market value of nwc and fixed assets = 115000+8000000 = $8115000